OpenStudy (perl):

Genius Question: Why is i^i^i^i^i^i^e = 1 where i = sqrt(-1)

OpenStudy (perl):

please show why this is equal to 1, you can use exponential polar form, for example i = exp(pi/2*i)

OpenStudy (anonymous):

Look at \[\large 2^{(3^4)}=2417851639229258349412352\\ (2^ 3)^4=4096 \] Where are your parentheses in your question?

OpenStudy (perl):

yes i know that

OpenStudy (perl):

i^i^i^i^i^i^e =i^(i^(i^(i^(i^(i^e))))

OpenStudy (perl):

exponentiation is right associative

OpenStudy (ikram002p):

i^i^i^i^i^i^e = 1 take ln for both e ln i^i^i^i^i^i =0 so ln i^i^i^i^i^i =0 then must be i^i^i^i^i^i =1 take ln i ln i^i^i^i^i =0 i^i^i^i^i =1 ln again i ln i^i^i^i =0 i^i^i^i =1 ln also i ln i^i^i =0 i^i^i =1 ln again i ln i^i =0 i^i =1 i ln i =0 i=1 its a contradiction so im not sure there is somthing wrong i think in my answer

OpenStudy (perl):

let me check that over, it looks wrong

OpenStudy (perl):

you took ln of both sides

OpenStudy (perl):